# Conversational Complexity for Assessing Risk in Large Language Models

### v1

@@ II.3 Conversational Complexity
where $\breve{C}=\langle u_{1},u_{2},...,u_{n}\rangle$ represents the sequence of user utterances in conversation C, and $K_{U}$ is the Kolmogorov complexity with U as the reference machine.
@@ V.1 Conversational Length, Conversational Complexity and Harm
Figure 4: Conversational Complexity against Conversational Length (in bits).
@@ V.2 Comparison of Model Types
Comparing across model types reveals several key insights. Models that incorporate red team data, such as RLHF and potentially Rejection Sampling, show less distinct separation between harmful and harmless conversations in terms of CC. This aligns with previous findings from [18] that these techniques result in less harmful model outputs overall. However, even with these advanced safety techniques, the distinction in CC between harmful and harmless conversations persists.
@@ VI Limitations and Potential
The Anthropic Red Teaming dataset, while valuable, presents its own challenges. Our tiered approach to categorizing harm, while necessary for analysis, may oversimplify the multifaceted nature of potential negative impacts from LLM outputs. A more nuanced, multi-dimensional framework for harm assessment could provide richer insights.
@@
Despite these limitations, our work presents significant potential for advancing LLM safety. We introduce a novel risk assessment framework based on Minimum Conversational Complexity (MCC), defined as the minimum Kolmogorov complexity of the user’s side of a conversation that elicits a specific output from an LLM. This approach allows us to quantify risk without relying on hard-to-estimate probabilities of user intentions and behaviors. Our risk function is based on a universal distribution of risk ([25]):
@@
Where $U$ is the user, $M$ is the model, ${\cal C}_{U,M}$ represents all possible conversations between the user and the model, and $\text{Harm}(C)$ encapsulates the potential harm of conversation $C$. This distribution weights simple, harmful conversations more heavily than complex ones, aligning with the intuition that easier-to-execute harmful interactions pose a greater risk. Given a sample of cases, instead of the full set ${\cal C}$, we can estimate this risk, as we do in the last column of Table 2. We also saw that by looking at the conversation with Minimum Conversational Complexity, we can also have a very good proxy of this (as shown in the next but last column of the table), because the Minimum Conversational Complexity dominates the sum in Eq. 1.
By focusing on interaction complexity rather than estimating specific user behavior probabilities, we offer a more tractable approach to risk assessment in AI systems. This framework has the potential to enhance red teaming methodologies by providing quantitative measures of conversation complexity and potential harm. It can be applied to estimate the autonomy of LLM agents in acquiring capabilities that lead to harm [33, 22], and help LLM developers prioritize their efforts in patching detected risks based on the complexity and potential harm of vulnerable interaction patterns.
@@ VII Acknowledgments
We utilized Anthropic’s Claude and OpenAI’s ChatGPT for editorial assistance during the preparation of this manuscript. These language models helped refine the paper’s language and structure.

### v2

@@ I Introduction
where M is the LLM being evaluated, U represents the user, C represents the conversation, and o represents the output.
@@
While ”conversational complexity” has been defined in various ways in the literature, our approach diverges significantly from prior conceptualizations. For example, [16] define conversational complexity in terms of how individuals cognitively differentiate and psychologically structure conversations, focusing on constructs like topic familiarity and enjoyment. Similarly [3] conceptualize complexity in terms of the miscalibration of learning expectations in conversations with strangers. These are unrelated redefinitions compared to our use of the term in the context of LLMs, where complexity is grounded in algorithmic information theory. Our approach aligns more closely with recent work [bergey2024uma] apples similar information-theoretic measures to human conversation. While they focus on natural human speech patterns, their method of quantifying conversational complexity provides a useful parallel to our LLM-based approach.
@@ II.3 Conversational Complexity
where $\breve{C}=\langle u_{1},u_{2},...,u_{n}\rangle$ represents the sequence of user utterances in conversation C, $K_{U}$ is the Kolmogorov complexity with U as the reference machine, and $h_{i}=\textlangle u_{1},m_{1},...,u_{i},m_{i}\textrangle$ represents the conversation history up to and including the i-th turn.
@@
In practice, computing $MCC(o)$ over all possible conversations is infeasible. Instead, we approximate MCC using carefully curated datasets designed to probe model behaviors, particularly those aimed at eliciting potentially harmful or undesired outputs. It’s important to note that the choice of dataset can significantly impact the estimated MCC values.
@@ V.1 Conversational Length, Conversational Complexity and Harm
Figure 4: Conversational Complexity against Conversational Length (in bits). The Pearson correlation coefficient between CC and CL is 0.949.
@@ V.2 Comparison of Model Types
Comparing across model types reveals several key insights. Models that incorporate red team data, such as RLHF and potentially Rejection Sampling, show less distinct separation between harmful and harmless conversations in terms of CC. Safety techniques tend to produce heavier tails for both ’harmful’ (successful) and ’harmless’ (failed) harm attempts, causing a greater overlap between these distributions. This increased complexity across both categories suggests that adversarial strategies are likely becoming more sophisticated in response to improved safety measures. The heavier tails in ’harmless’ conversations likely represent complex evasion attempts that ultimately failed to bypass the model’s safeguards. However, the distinction in CC between harmful and harmless conversations persists.
@@ V.3 Power Law Analysis of Complexity Distributions
It is important to note that while our analysis suggests power-law-like behavior in the distributions of Conversational Length and Conversational Complexity, the range of our data on the horizontal axis does not span multiple orders of magnitude, which is typically desired for a definitive power law identification. This limitation is inherent to the nature of our dataset and the practical constraints of human-LLM interactions. Despite this constraint, the observed distributions exhibit characteristics consistent with power laws within the available range. We interpret these results as indicative of scale-free properties in the conversation structures, rather than as definitive proof of power law behavior.
@@ VI Limitations and Potential
The Anthropic Red Teaming dataset, while valuable, presents its own challenges. Our tiered approach to categorizing harm, while necessary for analysis, may oversimplify the multifaceted nature of potential negative impacts from LLM outputs.
@@
Despite these limitations, our work presents significant potential for advancing LLM safety. We introduce a novel risk assessment framework based on Minimum Conversational Complexity (MCC), defined as the minimum Kolmogorov complexity of the user’s side of a conversation that elicits a specific output from an LLM. This approach allows us to quantify risk without relying on hard-to-estimate probabilities of user intentions and behaviors.
Our risk function is based on a universal distribution of risk ([26]):
@@
Where $U$ is the user, $M$ is the model, ${\cal C}_{U,M}$ represents all possible conversations between the user and the model, and $\text{Harm}(C)$ encapsulates the potential harm of conversation $C$. This distribution weights simple, harmful conversations more heavily than complex ones, aligning with the intuition that easier-to-execute harmful interactions pose a greater risk. Furthermore, due to the dominance property of Levin’s Universal Distribution, the Universal Risk Function serves as an upper bound on the overall risk, ensuring that our risk assessments remain conservative and robust against easily executable harmful interactions (see Appendix C).
Given a sample of cases, instead of the full set ${\cal C}$, we can estimate this risk, as we do in the last column of Table 2. We also saw that by looking at the conversation with Minimum Conversational Complexity, we can also have a very good proxy of this (as shown in the next but last column of the table), because the Minimum Conversational Complexity dominates the sum in Eq. 1.
By focusing on interaction complexity rather than estimating specific user behavior probabilities, we offer a more tractable approach to risk assessment in AI systems. Nevertheless, it’s important to acknowledge that this method has limitations due to its underlying assumptions about user input probabilities (see Appendix C).
This framework has the potential to enhance red teaming methodologies by providing quantitative measures of conversation complexity and potential harm. It can be applied to estimate the autonomy of LLM agents in acquiring capabilities that lead to harm [34, 23], and help LLM developers prioritize their efforts in patching detected risks based on the complexity and potential harm of vulnerable interaction patterns.
@@ VII Acknowledgments
We utilized Anthropic’s Claude and OpenAI’s ChatGPT for editorial assistance during the preparation of this manuscript. These language models helped refine the paper’s language and structure. We thank Miguel Ruiz Garcia for his valuable feedback on an earlier version of this manuscript.
@@ Appendix B Estimating Conversational Complexity using Compression
Appendix C Limitations of the Universal Risk Function
The Universal Risk Function assesses the risk associated with conversations between users and Large Language Models (LLMs) by weighting the potential harm of each conversation by the exponential of the negative Kolmogorov Complexity of the user’s input:
$\text{Risk}(U,M)=\sum_{C\in\mathcal{C}_{U,M}}2^{-K(\breve{C})}\cdot\text{Harm}(C),$ (4)
where $\mathcal{C}_{U,M}$ denotes the set of all possible conversations between user $U$ and model $M$, $\breve{C}$ represents the user’s input in conversation $C$, $K(\breve{C})$ is the Kolmogorov Complexity of $\breve{C}$, and $\text{Harm}(C)$ quantifies the potential harm of the conversation.
A fundamental assumption in this framework is that the probability of a user input $\breve{C}$ occurring is proportional to $2^{-K(\breve{C})}$, implying an exponential decay of input probabilities with increasing Kolmogorov Complexity:
$P(\breve{C})\propto 2^{-K(\breve{C})}.$ (5)
However, this assumption may not hold in practice, as real user inputs may not exhibit an exponential decrease in probability with increasing complexity due to multiple factors. Additionally, users may deliberately construct complex inputs to test the capabilities of LLMs or attempt to circumvent safety measures. By assigning lower probabilities to complex user inputs, the Universal Risk Function may underestimate the risk associated with harmful outputs elicited by such inputs. Conversely, it may overestimate the risk associated with simpler inputs.
Despite these limitations, the Universal Risk Function serves as an upper bound on the overall risk due to its foundational reliance on Levin’s Universal Distribution. Specifically, for any computable distribution $P(\breve{C})$, there exists a constant $c\geq 1$ such that:
$P(\breve{C})\leq c\cdot 2^{-K(\breve{C})}.$ (6)
This inequality ensures that the actual expected risk, defined as $\sum_{C\in\mathcal{C}_{U,M}}P(\breve{C})\cdot\text{Harm}(C)$, does not exceed $c$ times the Universal Risk Function $\text{Risk}(U,M)$. Consequently, the Universal Risk Function provides a conservative overestimation of the true risk, capturing worst-case scenarios and guiding the development of safety measures that are robust against inputs of minimal complexity.